# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 3.4.2 Bias NLI
Figure 27 provides a network visualization of the Top-15 PMI pairs, where nodes represent words and edges indicate high-PMI group–trait associations. Blue nodes represent group terms; green nodes represent trait terms. The graph reveals clusters of gender-based emotional traits (him–shy, women–shy), racial-socioeconomic links (black–unsuccessful, african–poverty), and age-personality pairings (young–confident, old–mean). Importantly, not all associations are overtly negative; terms like (european, kind) and (adult, confident) may reflect prescriptive stereotypes or privilege biases that still distort model learning.
@@
In sum, the presence of both positive and negative associations confirms that BiasNLI reflects not only harmful tropes but also normative assumptions about identity and capability. Since many of these associations emerge in hypothesis sentences—where models are known to rely heavily on surface cues—the leakage poses a substantial threat to the integrity of inference tasks. Debiasing strategies for template-based NLI construction should therefore consider lexical correlation audits as part of the generation pipeline.
Differential Metric Bias. Differential Matrics Bias analysis is a common approach in fairness research, typically used to compare the average values of various metrics (such as sentiment, toxicity, regard, or gender polarity) across different demographic groups within a dataset. This method relies on the presence of explicit, gold-standard demographic annotations—such as gender, race, or religion labels—for each example, enabling robust group-wise comparisons. However, the BiasNLI dataset does not provide such explicit demographic labels. Instead, any attempt to infer group membership must rely on heuristic keyword matching within the text (e.g., assigning a sentence to the “male” group if it contains words like “he” or “man”). This approach is inherently limited: it may miss implicit references, misclassify ambiguous cases, and cannot distinguish between subject and object mentions or account for non-binary identities. Furthermore, the presence of a demographic keyword in the text does not necessarily indicate the true group membership of the main entity or author, but merely that the group is mentioned. As a result, any group-based analysis in BiasNLI would reflect only the distribution of model scores for sentences mentioning certain demographic terms, rather than systematic differences across real-world groups. Given these limitations, we do not report Differential Matrics Bias results for BiasNLI, as such analysis would be unreliable and potentially misleading.
3.5 Conclusion on Counterfactual Datasets
Counterfactual datasets have played a foundational role in the fairness evaluation of language models by enabling controlled experiments that isolate the effects of identity attributes—such as gender, race, or occupation—through minimally varied sentence pairs. Early benchmarks like WinoBias and Winogender introduced template-based designs focused on binary gender stereotypes, while later datasets such as GAP, EEC, and CrowS-Pairs expanded demographic coverage and introduced more naturalistic or sentiment-oriented contexts. This evolution has continued with datasets like StereoSet, HolisticBias, and WinoQueer, which address underrepresented identities, incorporate participatory design, or explore sociolinguistic variation. However, challenges remain. Template-based corpora offer experimental control at the expense of realism, while naturalistic or crowdsourced datasets provide ecological validity but suffer from confounds and potential annotation bias. In addition, the predominance of English and Western-centric demographic framing limits cross-cultural generalization. Researchers must therefore navigate trade-offs between control, diversity, and representativeness, critically assessing dataset construction, target biases, and known limitations. Ultimately, while counterfactual datasets offer indispensable tools for probing specific fairness concerns, they are best employed alongside complementary evaluation methods and with an understanding of their methodological boundaries to ensure robust, inclusive, and contextually grounded fairness assessments.
4 Prompt-based Datasets
Prompt-based datasets represent a significant category of resources for evaluating fairness in language models (LMs). Unlike counterfactual datasets that rely on minimal-pair comparisons, prompt-based datasets assess fairness by analysing the content generated by an LM in response to an initial prompt or question. This mirrors real-world LM applications (text completion, dialogue, QA). Evaluation typically examines generated text for bias manifestations such as toxicity, stereotype reinforcement, or differential treatment triggered by demographic characteristics present or implied in the prompt. As defined in Section 3.1, prompt-based datasets are further subdivided into Sentence Completion and Question-Answering (QA) formats, reflecting distinct modes of LM interaction and evaluative focus.

### v2

@@ 3.4.2 Bias NLI
Figure 27 provides a network visualization of the top 15 PMI pairs, where nodes represent words and the edges indicate high-PMI group–trait associations. Blue nodes represent group terms; green nodes represent trait terms. The graph reveals clusters of gender-based emotional traits (him–shy, women–shy), racial-socioeconomic links (black–unsuccessful, african–poverty), and age-personality pairings (young–confident, old–mean). Importantly, not all associations are overtly negative; terms like (european, kind) and (adult, confident) may reflect prescriptive stereotypes or privilege biases that still distort model learning.
@@
In sum, the presence of positive and negative associations confirms that BiasNLI reflects not only harmful tropes but also normative assumptions about identity and capability. Because many of these associations emerge in hypothesis sentences, where models are known to rely heavily on surface cues, this leakage poses a substantial threat to the integrity of inference tasks. Consequently, debiasing strategies for template-based NLI construction should include lexical-correlation audits in the generation pipeline.
3.5 Multiple-Choice Question Answering Bias Datasets
This subsection introduces datasets that evaluate bias in question-answering (QA) systems, specifically in the context of multiple-choice questions and underspecified question formats. BBQ and UnQover test whether QA models exhibit biased behavior in choosing answers based on demographic cues embedded in the questions. For instance, BBQ provides multiple-choice questions that include stereotypical and non-stereotypical answers, measuring whether the model tends to select stereotypical answers when demographic cues like gender or race are involved. UnQover, on the other hand, focuses on questions that are intentionally underspecified, evaluating how models deal with ambiguity and whether their answers are influenced by biases related to gender, religion, or ethnicity.
3.5.1 BBQ (Bias Benchmark for QA)
Dataset Taxonomy. BBQ is classified as a constrained-form evaluation dataset within the multiple-choice question answering bias category. It employs a dual-context structure with both ambiguous contexts (where insufficient information is provided) and disambiguated contexts (with clear cues), allowing explicit measurement of bias when models must choose between socially charged inferences and uncertainty. Each question is paired with three answer options, one of which typically conveys a stereotype-aligned response, enabling evaluation of whether QA models exhibit biased behavior in selecting answers based on demographic cues embedded in the questions.
In terms of source, BBQ employs template-based construction combined with crowdsourced validation. The dataset was created manually by researchers using templates that reflect documented social stereotypes, then subsequently validated through crowdsourcing. This hybrid approach balances experimental control with human validation while potentially introducing inconsistencies in quality and demographic biases from the annotators themselves.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on U.S. English. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, BBQ targets multiple demographic characteristic biases across nine dimensions: age, disability status, gender identity, nationality, physical appearance, race/ethnicity, religion, sexual orientation, and socioeconomic status. The dataset purposefully targets demographic biases across these dimensions, providing a comprehensive coverage of protected attributes.
Finally, BBQ demonstrates high accessibility as a publicly available dataset via GitHub. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
Intrinsic Characteristics. The dataset consists of over 58,000 examples derived from 325 unique templates. Each instance contains a context (either ambiguous or disambiguated), a question and three candidate answers, including an “Unknown” option that is phrased in ten distinct ways to minimize lexical bias. The model predictions are evaluated using both accuracy and a dedicated Bias Score, which measures preference for stereotype-aligned answers over the “Unknown” choice in ambiguous contexts. This score ranges from 1-1 (anti-stereotype bias) to +1+1 (pro-stereotype bias). Validation by Mechanical Turk workers ensures alignment between gold labels and common-sense reasoning.
Bias Analysis.
Representativeness Bias. We assess representativeness by computing, for each protected-attribute axis $A$, the divergence between the empirical category distribution in BBQ, $P_{D}(A)$, and the target population distribution from BLS 2024, $P_{P}(A)$, using the KL definition in Equation 1. Concretely, for each axis, we estimate category frequencies from BBQ (blue bars in Figure 28) and compare them to the corresponding BLS 2024 proportions (red bars), then aggregate category-level contributions into a single axis-wise KL score (visualized in Figure 29). This procedure yields a spectrum of divergences: Gender is effectively aligned with the population baseline (KL =0.0024=0.0024), Religion remains low (0.1516), and Nationality (0.4800) and Race/Ethnicity (0.4968) are moderate. In contrast, Age Group (0.6622), Socioeconomic Class (0.8093), and Physical Appearance (0.8661) show substantial skew, while the two largest divergences arise for Sexual Orientation (1.5302) and Disability Status (1.5418). The underlying panels in Figure 28 clarify the drivers of these values: some population segments are missing or nearly missing in BBQ (e.g., abled and middle-income groups), and others are strongly amplified (e.g., LGBTQ+ and disabled). Such zero-zero or near-zero mass categories in $P_{D}$ inflate the corresponding terms in Equation 1, producing the largest KL scores. These distributional distortions have two methodological consequences for fairness evaluation: first, they can overemphasize bias behavior on amplified minorities while limiting coverage and statistical power for majority groups; second, they restrict intersectional analysis by leaving certain demographic intersections sparsely represented or absent. Accordingly, BBQ is most informative when paired with complementary resources that better approximate real-world population structure for the axes with high divergence.
Figure 28: Per-axis composition comparison for BBQ vs. BLS 2024. Each panel shows category proportions for BBQ (blue) and the BLS baseline (red) along one demographic axis. These empirical proportions $P_{D}(A)$ and $P_{P}(A)$ are the inputs to the axis-wise KL computation in Equation 1. Notable gaps include the absence or near-absence of some groups (e.g., abled, middle-income), and amplification of others (e.g., LGBTQ+, disabled).
Figure 29: Representativeness bias (KL) of BBQ relative to BLS 2024 by axis. Lower values indicate better alignment with the population baseline. The ordering highlights minimal divergence for Gender (0.0024) and Religion (0.1516), moderate values for Nationality (0.4800) and Race/Ethnicity (0.4968), and the largest divergences for Sexual Orientation (1.5302) and Disability Status (1.5418).
Annotation Bias. BBQ’s construction aims to avoid annotation bias by ensuring that templates instantiated across demographic groups share the same gold label, and ambiguous contexts consistently require the “Unknown” answer. This design principle makes the intended labeling function group-invariant, such that $\hat{B}_{\text{ann}}$ should approach zero. However, when introducing an automated labeling function $g_{\theta}$ via the Perspective API, evidence of annotation bias emerges, particularly for the Age attribute, which we analyzed as a representative case. To assess potential bias in the BBQ dataset from a toxicity perspective, we conducted a fine-grained linguistic analysis targeting the Age category. The analysis leveraged the Perspective API to score QA pairs on five toxicity-related dimensions: TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, and THREAT. Our focus was to determine whether answers referring to stereotyped age groups (e.g., “old” individuals) systematically exhibited higher toxicity than their “non-old” or “unknown” counterparts. To preserve contextual meaning, each answer was paired with its corresponding question to construct complete QA utterances (e.g., “Who was not comfortable using the phone? The grandfather.”), ensuring the Perspective API operated on linguistically complete units. This approach allowed us to simulate real-world toxicity triggers, capturing the implicit social framing embedded in QA construction.
Figure 30 displays boxplots for five toxicity-related metrics—TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, and THREAT—grouped by stereotype label. Although absolute toxicity scores remain low across all categories, a consistent pattern emerges: answers labeled as “Old” exhibit slightly higher medians and upper quartiles across multiple dimensions compared to their “Non-Old” and “Unknown” counterparts. This suggests that stereotype-aligned completions, such as “The grandfather”, are more likely to carry marginally elevated toxicity, even when no explicit slurs or offensive language are present. Notably, “Unknown” answers (e.g., “Can’t be determined”) consistently produce the lowest toxicity scores, reinforcing their role as a neutral linguistic baseline. These subtle but persistent disparities indicate that even in a carefully curated dataset like BBQ, implicit bias can manifest through contextual framing rather than overt lexical content. In our framework, this constitutes a non-zero empirical annotation bias as defined in Equation 2, arising not from BBQ’s human gold labels but from $g_{\theta}$’s priors over age-marked lexemes and the contextual frames in which they appear. Our analysis of the Age attribute demonstrates how auxiliary scorers, if used to audit or augment datasets like BBQ, can inadvertently reintroduce annotation bias, underscoring the need for calibration techniques such as template-level normalization, attribute-swapped evaluation, and paraphrase balancing to mitigate such effects.
Stereotype Leakage. Unlike naturalistic corpora where stereotype leakage manifests through implicit co-occurrence between demographic markers and trait terms, the BBQ dataset does not provide a suitable structure for such analysis. Each instance in BBQ is composed of a context, a question, and three answer options, one of which is explicitly stereotype-aligned. This template-based construction encodes stereotypes at the level of answer alternatives rather than allowing them to emerge through distributional patterns in free text. As a result, pointwise mutual information (PMI) and mutual information (MI), the information-theoretic estimators typically used to measure stereotype leakage by aggregating co-occurrence statistics (Equation 1011), cannot be applied meaningfully to BBQ. Instead, the presence of stereotypes in BBQ is a designed feature of the dataset, purposefully embedded to evaluate whether models rely on stereotype-consistent inferences when contexts are ambiguous. Thus, while BBQ is highly effective as a diagnostic resource for probing stereotype-driven reasoning in question answering, it does not support corpus-level leakage estimation within our framework.
3.5.2 UnQover
Description. UnQover is a diagnostic dataset and evaluation framework designed to examine stereotyping biases in both Question Answering (QA) systems and Masked Language Models (MLMs). Introduced by Li et al. [29], it generates underspecified questions using fewer than 30 manually designed templates. These templates combine contextual phrases with questions such as “Who [attribute/activity]?”, where multiple demographic identities (e.g., gender, nationality, religion) could be equally plausible. The ambiguity is intentional, as by omitting key disambiguating information, the dataset aims to surface latent stereotypes encoded in models’ preferences or scoring patterns.
Dataset Taxonomy. UnQover is classified as a constrained-form evaluation dataset within the multiple-choice question answering bias category. It generates underspecified questions using fewer than 30 manually designed templates that combine contextual phrases with questions such as "Who [attribute/activity]?", where multiple demographic identities could be equally plausible. The intentional ambiguity aims to surface latent stereotypes encoded in models’ preferences or scoring patterns, enabling evaluation of both QA systems and Masked Language Models through systematic bias detection in ambiguous scenarios.
In terms of source, UnQover employs template-based construction with a generation strategy using predefined slots populated by identity groups and stereotypically associated attributes or activities. This approach provides precise control over sentence variability and demographic terms, allowing for targeted fairness evaluations while maintaining experimental rigor through systematic template generation.
The dataset exhibits monolingual linguistic coverage, focusing primarily on English. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, UnQover targets multiple demographic characteristic biases, specifically those related to gender, ethnicity, nationality, and religion. The dataset focuses on identity-based stereotyping in language models, especially where model behavior may implicitly reflect social biases despite neutral inputs, providing comprehensive coverage of protected attributes.
Finally, UnQover demonstrates high accessibility as a publicly available dataset and evaluation toolkit on GitHub. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
Intrinsic Characteristics. UnQover is designed for large-scale generation, producing evaluation files that can span several gigabytes and potentially include millions of instances. Each instance comprises a short context followed by an ambiguous question. For MLMs, corresponding templates replace answer spans with <mask> tokens. The subjects in each prompt pair are drawn from contrasting demographic categories (e.g., male vs. female names), and attributes are carefully chosen for their stereotypical connotations. The evaluation procedure is score-based rather than purely answer-based: key metrics include the Subject Bias score (C), Position Error (Δ), and Attribute Error (ε), which capture whether a model systematically favors one group in ambiguous scenarios.
Strengths and Limitations. UnQover’s strengths lie in its systematic design, fine-grained metrics, and high scalability. Its use of underspecified prompts offers a clean testbed for bias attribution, and its compatibility with both QA and MLM architectures broadens applicability. Furthermore, the public release of generation and scoring code enhances transparency and reproducibility. However, the approach is not without limitations. The reliance on templated inputs may fail to capture the richness and ambiguity of natural language, and deliberate selection of stereotype-triggering attributes may raise ethical concerns. Additionally, as with other template-based datasets, such as BBQ, its English-centric focus and simplified structure may limit generalizability across languages or nuanced contexts.
Bias Analysis.
Representativeness Bias. Unlike BBQ, which is template-based yet designed for broad demographic coverage across 11 protected attributes and therefore invites comparisons between dataset composition and societal demographics, UnQover is a synthetically generated, contrastive benchmark that does not aspire to reflect population-level frequencies. Its instances are produced by slot-filling a small set of templates that deliberately balance identity pairs (e.g., male vs. female, Christian vs. Muslim), so applying our KL-divergence criterion for representativeness bias (Equation 1) is not strictly meaningful. Nevertheless, UnQover may exhibit a design-level representativeness limitation: the specific identities and stereotype-triggering attributes selected for inclusion (and omission) constrain the evaluation space, privileging widely recognized groups and commonly discussed attributes while under-covering others. Thus, while BBQ’s internal composition supports a quantitative representativeness analysis though $D_{\mathrm{KL}}(P_{D}(A)\,\|\,P_{P}(A))$, UnQover is better assessed qualitatively for coverage scope—i.e., whether its template and attribute choices yield a sufficiently diverse identity coverage for the claims an experiment seeks to make.
Annotation Bias. Annotation bias is not directly applicable to the UnQover dataset. By construction, UnQover does not include human- or model-assigned gold labels; rather, it consists of templated, underspecified questions designed to elicit model predictions in ambiguous contexts. The absence of reference labels means that systematic discrepancies in labeling across demographic groups, our formal definition of annotation bias, cannot arise. Thus, while annotation bias is a critical dimension for many labeled fairness datasets, it is not a meaningful analytic lens for UnQover, whose diagnostic power derives precisely from omitting labels and observing model preferences under ambiguity.
Stereotype Leakage. Unlike sentence-based corpora such as WinoBias or BOLD, the UnQover dataset is not composed of declarative linguistic items, but rather underspecified question–answer pairs generated from manually designed templates. Each instance presents a short context (optional), an ambiguous question (e.g., “Who is good at math?”), and multiple candidate answers that correspond to demographic identities. Therefore, the bias signal in UnQover arises from the model’s resolution of ambiguity, i.e. which candidate answer it favours systematically, rather than from the naturalistic co-occurrence patterns of group words and trait descriptors.
Because our stereotype leakage framework relies on information-theoretic statistics (PMI and MI) over co-occurrence counts between identity markers $W_{G}$ and trait words $W_{T}$, it cannot be directly applied to UnQover instances. The dataset does not produce co-occurrence events in declarative text but instead evaluates differential model scoring across logically equivalent candidate answers. As a result, while UnQover provides a robust diagnostic for stereotyping in QA and MLM settings, its design precludes the computation of stereotype leakage as defined in our framework. Future work might extend leakage estimators to decision-based formats, for example, by operationalizing systematic answer preference as an analogue to corpus-level association.
3.6 Information Retrieval Bias Datasets
This subsection covers datasets designed to assess bias in information retrieval systems. These datasets evaluate whether search engines or document retrieval models favor certain demographics, such as gender, race, or occupation, when ranking or selecting results. Grep-BiasIR is the primary dataset in this category, testing whether search queries with embedded identity-related terms result in biased or stereotypical retrieval outcomes. It helps evaluate if retrieval systems exhibit preferences based on demographic attributes, such as favoring stereotypical gender or occupation associations in search results.
Dataset Taxonomy. Grep-BiasIR is classified as a constrained-form evaluation dataset within the information retrieval bias category. It consists of gender-neutral yet bias-sensitive queries, each paired with document sets that are systematically manipulated to reflect different gender indicators through controlled substitution of gender-indicating terms such as pronouns or role nouns. The dataset evaluates whether information retrieval systems preferentially retrieve or rank documents associated with specific genders, even when the query itself does not imply any gendered intent, providing a framework for assessing representation bias in ranking systems.
In terms of source, Grep-BiasIR employs a hybrid approach combining natural text sources with template-based modification. Queries are manually crafted along gender stereotype dimensions, and associated documents are drawn from natural sources such as web search results, then modified via word substitution to generate gendered variants. This approach provides both ecological realism and experimental control, while potentially introducing construction artifacts from the modification process.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on English. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, Grep-BiasIR focuses primarily on gender bias, representing a demographic characteristic bias with a focus on gender representation. The dataset covers seven stereotype dimensions: Appearance, Child Care, Cognitive Capabilities, Domestic Work, Career, Physical Capabilities, and Sex & Relationship, providing comprehensive coverage of gender-related bias contexts.
Finally, Grep-BiasIR demonstrates high accessibility as a publicly available dataset though GitHub. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
Intrinsic Characteristics. The dataset contains 118 queries and 708 documents, with each query linked to a relevant and a non-relevant document in all three gendered forms. All queries are gender-neutral by construction, ensuring that any bias observed in retrieval or ranking arises from the system’s treatment of the content, not the query phrasing. The dataset underwent expert auditing by post-doctoral researchers, who annotated the documents with expected stereotype directions and verified quality. Evaluation typically involves analyzing how an IR system ranks gendered versions of the same document, revealing possible systematic preferences. In addition to automated relevance scoring, the dataset also supports user studies for evaluating perceived relevance across gendered variants.
Domain Focus and Significance. Grep-BiasIR is situated in the domain of IR fairness and focuses on gender representation bias across seven stereotype dimensions: Appearance, Child Care, Cognitive Capabilities, Domestic Work, Career, Physical Capabilities, and Sex & Relationship. Unlike many bias benchmarks that focus on toxic language or explicit stereotype association, Grep-BiasIR foregrounds representational disparity, the skew in how frequently systems surface or prioritize content aligned with particular identities in response to ostensibly neutral queries.
Strengths and Limitations. Grep-BiasIR provides a rare and rigorously constructed resource for studying representation bias in IR, featuring high-quality, minimally contrastive document sets that enable precise attribution of system behavior to gender indicators. Its methodological clarity and breadth of stereotype coverage make it a strong baseline for future research. However, its limitations include a relatively small scale and focus on binary gender identities, with no representation of non-binary individuals. Cultural specificity in the stereotype framing may also affect generalizability. Moreover, while its IR design is well-suited for retrieval-based evaluation, applying it to modern LLMs may require reframing the task, such as assessing summaries generated for gendered versions of the same content. Nonetheless, the dataset’s emphasis on representation bias makes it a valuable template for adapting such methodologies to broader fairness evaluations beyond IR.
Bias Analysis.
Representativeness Bias. The design of Grep-BiasIR minimizes representativeness bias within its own scope by enforcing a balanced construction across gender categories: each query is paired with three document variants (male, female, and neutral), ensuring that $P_{D}(A)$ is symmetric with respect to binary gender indicators. Consequently, compared to an idealized target population $P_{P}(A)$ limited to male and female groups, the KL divergence (Equation 1) would approach zero, indicating low intra-dataset representativeness bias. However, when the broader target population is considered, including non-binary, transgender, and culturally diverse gender identities, the dataset reveals structural underrepresentation, as these groups are entirely absent. Moreover, the stereotype domains (Appearance, Child Care, Cognitive Capabilities, Domestic Work, Career, Physical Capabilities, and Sex & Relationship) represent only a narrow slice of the gender-related contexts that appear in real-world IR tasks. Thus, Grep-BiasIR should be interpreted as a resource that foregrounds representational bias in a controlled binary setting rather than one that comprehensively reflects demographic distributions in natural populations.
Annotation Bias. In the case of Grep-BiasIR, annotation bias can in principle be assessed within our framework by measuring the maximum expected divergence in label distributions across gendered variants. Formally, one could estimate $\hat{B}_{ann}$ as defined in Equation 2), where the labels $y$ correspond to binary relevance judgments (relevant vs. non-relevant). Since each query is paired with minimally contrastive male, female, and neutral documents, systematic differences such that $\mathbb{E}_{x\sim D_{male}}[y]\neq\mathbb{E}_{x\sim D_{female}}[y]$ would signal annotation bias, revealing a skew in how annotators judged semantically equivalent content across genders. In practice, however, Grep-BiasIR employs a deterministic construction pipeline where relevance labels are mirrored across gendered variants and verified by expert auditors. As a result, empirical estimates of $\hat{B}_{ann}$ converge to near zero, indicating that conventional annotation errors are largely eliminated. However, subtler forms of annotation bias may persist at the design level: the cultural framing of stereotype dimensions, and the assumption that lexical substitutions (e.g., “mother” vs. “father”) preserve identical relevance. These structural choices, while not directly captured by Equation 2, illustrate how annotation processes can embed bias beyond simple discrepancies in label distributions.
Stereotype Leakage. Stereotype leakage in the Grep-BiasIR dataset is fundamentally neutralized by its counterfactual construction. Each base document is deliberately edited into minimally contrastive variants, where gender-indicating terms (e.g., “mother” vs. “father”) are systematically substituted while all other lexical material is preserved. Under the co-occurrence-based estimators defined in Eqs. 10-11, this design ensures that any observed associations between gender markers $W_{G}$ and trait terms $W_{T}$ arise directly from the enforced substitutions rather than from organic co-occurrence patterns. As a result, pointwise PMI and corpus-level MI values collapse toward uniformity across gendered variants, making the measurement of stereotype leakage uninformative for this dataset. In other words, Grep-BiasIR does not embed spontaneous demographic–trait linkages in the way that naturally occurring corpora do; instead, its value lies in enabling controlled evaluations of representation bias through document ranking. Consequently, applying stereotype leakage analysis to Grep-BiasIR is not meaningful, since the dataset’s design precludes the very type of implicit group–trait associations that the leakage metric is intended to capture.
3.7 Conclusion on Constrained‑Form Evaluation Datasets
Constrained‑form evaluation datasets have become indispensable for fairness research because they transform abstract ethical questions into quantifiable decision tasks. Across coreference, sentence‑likelihood, classification, multiple‑choice, and information‑retrieval benchmarks, a common design principle emerges: models are forced to choose among fixed outputs, allowing disparities to be traced directly to identity cues rather than to generative idiosyncrasies. Early corpora such as WinoBias and EEC offered tight experimental control through templates, thereby facilitating clean estimates of gender‑occupation bias. Subsequent resources, including CrowS‑Pairs, HolisticBias, and BBQ, broadened the demographic scope, introduced crowd‑authored or participatory sentences, and tested higher‑level reasoning, signalling a shift toward ecological validity.
However, three systematic bias dimensions persist. Representativeness bias remains pronounced: most datasets are English‑centric, Western‑framed, and skewed toward binary gender constructs, limiting conclusions about global language communities. Annotation bias also endures; crowd judgments in stereotype or sentiment tasks often reflect annotators’ own social positions, which can inflate or obscure measured gaps. Finally, stereotype leakage is evident whenever identity terms co‑occur disproportionately with evaluative adjectives or occupations, an artifact most visible in likelihood and sentiment benchmarks.
Trade‑offs are therefore unavoidable. Template collections maximise internal validity but risk sterile or unnatural phrasing; naturalistic or Reddit‑sourced corpora inject authenticity yet import confounds and annotator subjectivity; multi‑choice and ranking sets expose subtle preference biases but hinge on carefully engineered distractors and scoring rules. Researchers must align dataset selection with the target deployment context, triangulating evidence from multiple subcategories and, where possible, augmenting with multilingual or domain‑specific audits. Used judiciously and in concert with open‑ended evaluations, constrained‑form datasets provide a rigorous scaffold for diagnosing, comparing, and ultimately mitigating unfair decision patterns in modern language models.
4 Open-ended Evaluation Datasets
Figure 31: Overview of open-ended evaluation datasets discussed in Section 4.
Following our analysis of constrained-form evaluation datasets in Section 3, we now turn to open-ended evaluation datasets. Figure 31 provides an visual map of this family and the representative datasets considered here (BOLD, RealToxicityPrompts, HONEST, TrustGPT). Open-ended datasets are critical for assessing fairness in language models because they evaluate how models generate content in response to unstructured prompts, capturing biases that emerge in free-form generation across diverse contexts rather than in fixed outputs. They mirror real-world applications such as long-form text generation, dialogue, and open-domain question answering, and they surface risks, including stereotype reinforcement, toxicity, and disparate sentiment or regard. In what follows, we apply the analysis protocol from Section 2 to characterize their sources, linguistic coverage, bias typology, and accessibility, and we report dataset-level findings that inform practical dataset selection.